# History-Aware Neural Operator: Robust Data-Driven Constitutive Modeling of Path-Dependent Materials

### v1

@@ Abstract
This study presents an end-to-end learning framework for data-driven modeling of path-dependent inelastic materials using neural operators. The novel framework is built on the premise that the irreversible evolution of material responses, governed by hidden dynamics, can be inferred from observable data. We develop the History-Aware Neural Operator (HANO), an autoregressive model that predicts path-dependent material responses from short segments of recent strain–stress history without relying on hidden state variables, thereby overcoming the self-consistency issues commonly encountered in recurrent neural network (RNN)-based models. Built on a Fourier-based neural operator backbone, HANO enables discretization-invariant learning. To further enhance its ability to capture both global loading patterns and critical local path dependencies, we embed a hierarchical self-attention mechanism that facilitates multiscale feature extraction. Beyond ensuring self-consistency, HANO mitigates sensitivity to initial hidden states—a commonly overlooked issue that can lead to instability in recurrent models when applied to generalized loading paths. By modeling stress–strain evolution as a continuous operator rather than relying on fixed input–output mappings, HANO naturally accommodates varying path discretizations and exhibits robust performance under complex conditions, including irregular sampling, multi-cycle loading, noisy data, and pre-stressed states. We evaluate HANO on two benchmark problems: elastoplasticity with hardening and progressive anisotropic damage in brittle solids. Results show that HANO consistently outperforms baseline models in predictive accuracy, generalization, and robustness. With its demonstrated capabilities and discretization-invariant design, HANO provides an effective and flexible data-driven surrogate for simulating a broad class of inelastic materials and is well-suited for integration with classical numerical solvers.
@@
Data-driven modeling, Inelastic materials , Fourier neural operator, Path-dependency, Inelasticity , Operator learning, Thermodynamic consistency , Recurrent neural networks
@@ 1 Introduction
Constitutive modeling of history-dependent materials presents fundamental challenges in computational solid mechanics. The traditional phenomenological approach requires constructing mathematical models that describe material behavior through internal variables and evolution equations [46, 16]. These models often need extensive calibration and parameter identification, particularly challenging for materials exhibiting complex path-dependent responses such as plasticity, damage, or their combinations. These challenges have motivated a paradigm shift toward data-driven methods, which aim to learn material behavior directly from experimental or simulation data without imposing strong assumptions about the functional form of the constitutive relationships.
@@ 1.3 Proposed methodology, major contributions, and paper organization
To address both discretization sensitivity and initial-state dependence in modeling path-dependent inelastic materials, we propose a novel end-to-end data-driven constitutive modeling framework based on recent advances in neural operator learning [39, 36, 43], termed the history-aware neural operator (HANO). HANO builds on the Fourier neural operator (FNO)  [39] and adopts an autoregressive formulation inspired by Neural ODEs. By leveraging fast Fourier transforms, the embedded FNO component can capture global and long-range interactions in the frequency domain, enabling the model to effectively handle multi-scale phenomena and loading histories sampled at varying resolutions. Through its autoregressive structure, HANO infers future stresses directly from short segments of observable strain–stress history, eliminating the need for hidden state variables or integration from an undeformed reference configuration. This design mitigates self-consistency issues and reduces sensitivity to assumed initial conditions. The FNO backbone is further enhanced with a U-shaped architecture [52] and attention mechanisms [59], allowing the model to capture both global patterns (via spectral representations) and localized path dependencies (via convolutional paths and attention). As a result, the proposed HANO approach delivers accurate, discretization-invariant stress predictions and can be robustly applied to scenarios involving partial or irregularly sampled loading histories.
@@ 2.2 Data-driven constitutive learning for path-dependent materials
where $\bm{\varepsilon}(t)$ and $\bm{\sigma}(t)$ denote the strain and stress at current time $t$, $\{\bm{\varepsilon}(\tau),\bm{\sigma}(\tau)\}_{0\leq\tau\leq t}$ represents the prior strain–stress history, and $\mathcal{E}$ represents the evolution operator that maps the historical loading path into a latent state $\mathbf{z}(t)$ encoding the material memory. The goal is to predict the future stress state $\bm{\sigma}(t+\Delta t)$ based on the given future strain state $\bm{\varepsilon}(t+\Delta t)$ and the current internal states $\mathbf{z}(t)$.
@@
This mapping in Eq. (12) provides the mathematical foundation for many existing data-driven constitutive models, which aim to learn the underlying path-dependent material behavior from the recent history of observable states. However, how to accurately represent the evolution operator $\mathcal{E}$ in data-driven modeling remains an open challenge, leading to a variety of approaches.
@@ 2.3 Recurrent Neural Networks (RNNs)
Formally, forward propagation within a simple (tanh‑activated) RNN cell is expressed: For $n=1,...,N$,
@@ 2.4 Pitfalls of RNN-based constitutive modeling
[…]
@@ 3.1 Neural operators
where $\mathbf{a}(x)\in\mathcal{A}$ and $\mathbf{u}(x)\in\mathcal{U}$ are the input and output functions, respectively, with $x\in\mathcal{D}$ representing generalized coordinates defined in the parameter domain $D$. In Eq. (15), $\mathcal{P}$ and $\mathcal{Q}$ are lifting and projection operators (both typically implemented as multilayer perceptrons). Also, $\mathcal{L}_{\ell}=\rho\bigl{(}\mathcal{W}_{\ell}+\mathcal{K}_{\ell}\bigr{)}$ represents the operator of the $\ell$th hidden layer, i.e., $\mathcal{L}_{\ell}:\mathcal{V}\to\mathcal{V}$. In each layer, $\rho$ denotes a point-wise non-linearity activation function, $\mathcal{W}_{\ell}$ is a local linear operator, and $\mathcal{K}_{\ell}$ is an integral kernel operator.
@@ 3.1.1 Fourier neural operator
where $R_{\ell}(\mathbf{k})$ is specified only up to a cutoff frequency, providing spectral regularisation and a finite computational bandwidth. The layer output is then obtained by combining this global convolution with a point-wise linear map and a non-linearity. In summary, the operation of the $\ell$th hidden layer is,
@@ 3.3 History-aware neural operator and its implementation
As shown in Eq. (20), HANO extends the standard neural operator framework by interleaving multiple standard Fourier layers with AEUF layers. Specifically, it leverages $\mathcal{L}^{\text{AEUF}}$ layers to further refine the latent representations produced by the preceding Fourier layers, through coupling the global spectral context with attention-guided, multiscale detail recovery, as discussed in Section 3.2. The embedded self-attention module dynamically re-weights the strain–stress history, allowing the model to selectively focus on the most relevant segments for next-step prediction. Within each $\mathcal{L}^{\text{AEUF}}$ block, a U-Net pathway recovers high-frequency and localized features that may otherwise be attenuated by the truncated Fourier basis, while a parallel Fourier branch retains coarse-scale global trends. This combination of U-Net, Fourier, and linear branches enables the AEUF layer to effectively fuse low- and high-frequency information in a single pass. By stacking multiple AEUF layers, HANO progressively enhances its capacity to model both long-range dependencies and localized transitions, leading to substantial improvements in path-dependent stress prediction, as demonstrated in Section 5.3.
@@ 3.3.1 Data structure of HANO for constitutive modeling
Remark 3.3 (Hidden variable independence): Different from many existing data-driven approaches that rely on reconstructing (physical or nominal) hidden variables encode loading history information [20, 42, 30], HANO directly leverages available strain–stress histories as input. This end-to-end formulation allows HANO to operate as a hidden-state-free model while capable of capturing the underlying dynamics of internal variable (IVS) evolution without explicitly modeling them. By carrying history through observable data rather than latent states, HANO achieves consistent and robust predictions from any arbitrary point along the loading path. This design naturally supports autoregressive forecasting, as further demonstrated in 4.2, where the model accurately predicts path-dependent material behavior even when initialized mid-trajectory.
@@ 3.3.2 Model prediction & loss function
where $n$ denotes the current step, and $k$ denotes the chosen window length. For multi‑step forecasting, we iterate (22) recursively: after predicting $\hat{\bm{\sigma}}_{n+1}$, we update the strain $\bm{\varepsilon}_{n+1}=\bm{\varepsilon}_{n}+\Delta\bm{\varepsilon}_{n+1}$, slide the window forward by discarding the oldest pair $(\bm{\varepsilon}_{n-k+1},\bm{\sigma}_{n-k+1})$ and appending the newly formed pair $(\bm{\varepsilon}_{n+1},\hat{\bm{\sigma}}_{n+1})$, and then supply the next increment $\Delta\bm{\varepsilon}_{n+2}$. In short, we set $n\rightarrow n+1$ in Eq. (22) in the next step. Repeating this procedure until the desired prediction length yields the complete stress trajectory.
@@ 3.3.3 Training scheme
This training approach serves dual practical purposes. In early epochs, we keep inputs nearly clean, allowing the network to learn the fundamental stress-strain relationship without interference. As training progresses, we increase noise levels while simultaneously implementing scheduled sampling to phase out teacher forcing. By coupling scheduled sampling with noise injection, the training scheme gradually exposes the network to similar imperfect inputs it will create during inference. As teacher forcing is phased out and controlled noises are added, the model learns to attenuate its own prediction drift while remaining resilient to measurement noise. This joint strategy yields materially more reliable long-horizon stress forecasts, because the network is conditioned to operate under realistic inference conditions well.
@@ 4.1 Data preparation & problem setting
The training and testing datasets used in this verification example are derived from the one-dimensional elastoplastic model with kinematic hardening, characterized by the Young’s modulus $E=200\,\text{GPa}$, the yield stress $\sigma_{y}=0.2\,\text{GPa}$, the hardening modulus $H=20\,\text{GPa}$. Details of the constitutive model and its implementation are provided in A. Starting from a stress‐free state (referred to as the undeformed configuration), we generate 1,000 distinct loading histories, each comprising two complete loading–unloading cycles and uniformly sampled with 100 increments (i.e., $T_{j}\equiv 100,\forall j$). The loading and unloading strain amplitudes $e_{\text{load}}$ and $e_{\text{unload}}$ are randomly sampled from the interval [0.008, 0.015]0.0080.015[0.008,\,0.015][ 0.008 , 0.015 ] and [0.003, 0.007]0.0030.007[0.003,\,0.007][ 0.003 , 0.007 ], respectively. Unless otherwise stated, all stress‐related quantities in this section are expressed in gigapascals (GPa).
@@
The effectiveness of the proposed HANO model was evaluated through a benchmark test against two variants of RNN-based models: (1) the original RNN model from Gorji et al. [20], designated as RNN1, which exclusively uses $k$ strain history data as input, and (2) a modified RNN model, designated as RNN2, whose input layer is revised to receive the same concatenated strain–stress history window and forthcoming strain increment as the HANO model. All data-driven models (HANO and RNNs) are trained with the identical protocol described in Section 3.3.3.

### v2

@@ Abstract
This study presents an end-to-end learning framework for data-driven modeling of path-dependent inelastic materials using neural operators. The novel framework is built on the premise that the irreversible evolution of material responses, governed by hidden dynamics, can be inferred from observable data. We develop the History-Aware Neural Operator (HANO), an autoregressive model that predicts path-dependent material responses from short segments of recent strain–stress history without relying on hidden state variables, thereby overcoming the self-consistency issues commonly encountered in recurrent neural network (RNN)-based models. Built on a Fourier-based neural operator backbone, HANO enables discretization-invariant learning. To further enhance its ability to capture both global loading patterns and critical local path dependencies, we embed a hierarchical self-attention mechanism that facilitates multiscale feature extraction. Beyond ensuring self-consistency, HANO mitigates sensitivity to initial hidden states, a commonly overlooked issue that can lead to instability in recurrent models when applied to generalized loading paths. By modeling stress–strain evolution as a continuous operator rather than relying on fixed input–output mappings, HANO naturally accommodates varying path discretizations and exhibits robust performance under complex conditions, including irregular sampling, multi-cycle loading, noisy data, and pre-stressed states. We evaluate HANO on two benchmark problems: elastoplasticity with hardening and progressive anisotropic damage in brittle solids. Results show that HANO consistently outperforms baseline models in predictive accuracy, generalization, and robustness. With its demonstrated capabilities and discretization-invariant design, HANO provides an effective and flexible data-driven surrogate for simulating a broad class of inelastic materials.
@@
Data-driven modeling, Inelastic materials , Fourier neural operator, Path-dependency, Operator learning, Self-attention , Damage
@@ 1 Introduction
Constitutive modeling of history-dependent materials presents fundamental challenges in computational solid mechanics. Traditional phenomenological approaches rely on internal variables and evolution equations to describe material behavior [51, 17]. Developing such models demands extensive domain expertise for their formulation construction and often involves laborious parameter identification, which becomes particularly difficult for materials with complex path-dependent responses such as plasticity, damage, or their combinations. These challenges have motivated a paradigm shift toward data-driven methods, which aim to learn material behavior directly from experimental or simulation data without imposing strong assumptions about the functional form of the constitutive relationships. By replacing intricate model construction with learning from data, data-driven approaches offer greater flexibility in capturing complex behaviors that traditional methods often struggle to represent accurately.
@@ 1.2 Neural ODE-based approach
1.3 Thermodynamic Consistency in Data-Driven Constitutive Modeling
Another important class of data-driven constitutive modeling approaches focuses on embedding thermodynamic consistency constraints to ensure physical meaningfulness and improve generalization capabilities. These methods offer the significant advantage of guaranteeing that model predictions adhere to fundamental thermodynamic laws, providing theoretical rigor and enhanced reliability under limited training data conditions. Physics-informed constraints in these frameworks are typically implemented through three main strategies: (1) selecting network architectures that inherently satisfy constraints; (2) incorporating penalty terms in loss functions to prevent constraint violations; and (3) employing posterior rejection strategies that filter optimized models based on predefined acceptance/rejection criteria [17].
Thermodynamics-based Artificial Neural Networks (TANNs) proposed by Masi et al. [51, 49] represent a prominent methodology in this domain, which directly encodes thermodynamic principles into neural network architectures through dual potential formulations. Under isothermal conditions, this approach combines Helmholtz free energy density $F$ and dissipation rate potential $D$ using two independent neural networks: one for predicting internal variable increments and another for predicting Helmholtz free energy density. This architectural design ensures that stress updates strictly adhere to thermodynamic laws, particularly the second law of thermodynamics. Similarly, several studies [39, 64, 36, 58, 14] have also employed input convex neural networks to ensure material stability and thermodynamic consistency.
However, potential-based approaches face several practical challenges. The determination of appropriate structure and number of internal variables often remains empirical. Additionally, incorporating convexity constraints or dissipation potentials frequently requires complex optimization schemes, automatic differentiation, and numerical integration procedures. Moreover, these methods may exhibit sensitivity to temporal discretization due to reliance on explicit ODE solvers for state evolution [32], which can compromise the robustness objectives of operator learning approaches.
@@ 1.4 Proposed methodology, major contributions, and paper organization
To address both discretization sensitivity and initial-state dependence in modeling path-dependent inelastic materials, we propose a novel end-to-end data-driven constitutive modeling framework based on recent advances in neural operator learning [43, 40, 47], termed the history-aware neural operator (HANO). HANO builds on the Fourier neural operator (FNO)  [43] and adopts an autoregressive formulation inspired by Neural ODEs. By leveraging fast Fourier transforms, the embedded FNO component can capture global and long-range interactions in the frequency domain, enabling the model to effectively handle multi-scale phenomena and loading histories sampled at varying resolutions. To best of authors’ knowledge, this is the first application of neural operator architectures to history-dependent constitutive modeling. Through its autoregressive structure, HANO infers future stresses directly from short segments of observable strain–stress history, eliminating the need for hidden state variables or integration from an undeformed reference configuration. This design mitigates self-consistency issues and reduces sensitivity to assumed initial conditions. The FNO backbone is further enhanced with a U-shaped architecture [57] and attention mechanisms [66], allowing the model to capture both global patterns (via spectral representations) and localized path dependencies (via convolutional paths and attention). As a result, the proposed HANO approach delivers accurate, discretization-invariant stress predictions and can be robustly applied to scenarios involving partial or irregularly sampled loading histories.
@@ 2.2 Data-driven constitutive learning for path-dependent materials
where $\bm{\varepsilon}(t)$ and $\bm{\sigma}(t)$ denote the strain and stress at current time $t$, $\{\bm{\varepsilon}(\tau),\bm{\sigma}(\tau)\}_{0\leq\tau\leq t}$ represents the prior strain–stress history, $\Delta t$ represents an arbitrary time increment, and $\mathcal{G}$ represents the evolution operator that maps the historical loading path into a latent state $\mathbf{z}(t)$ encoding the material memory. The goal is to predict the future stress state $\bm{\sigma}(t+\Delta t)$ based on the given future strain state $\bm{\varepsilon}(t+\Delta t)$ and the current internal states $\mathbf{z}(t)$.
@@
This mapping in Eq. (12) provides the mathematical foundation for many existing data-driven constitutive models, which aim to learn the underlying path-dependent material behavior from the recent history of observable states. However, how to accurately represent the evolution operator $\mathcal{G}_{z}$ in data-driven modeling remains an open challenge, leading to a variety of approaches.
@@ 2.3 Recurrent Neural Networks (RNNs)
Formally, forward propagation within a simple (tanh‑activated) RNN cell is expressed: For n=1,,Nn=1,...,N,
@@ 2.4 Pitfalls of RNN-based constitutive modeling
    Another key limitation of conventional RNN-based constitutive models is their susceptibility on the initial hidden state. During training phase, the hidden vector $\mathbf{h}_{\,0}$ is typically initialized from a predefined reference configuration, most often in the undeformed zero-stress state, and is subsequently updated through the recursive relation in Eq. (13). Since $\mathbf{h}_{\,n}$ implicitly stores the entire loading history, any deviation from this training baseline during the inference/prediction phase can lead to significant distribution shift. For instance, when RNN predictions are employed in scenarios where the test specimen is already pre‑stressed or when the early portion of the strain-stress history is unavailable, the model must evolve from an incorrect $\mathbf{h}_{\,0}$. In such cases, the RNN is unable to fast-forward to the correct internal memory or reconstruct the missing increments, resulting in misalignment with the underlying evolution ODEs (Eq. (4)). As discussed in Section 4.2, the RNN model can result in large prediction errors when the model is forced to begin from mid‑sequence data rather than from an undeformed baseline.
@@ 3.1 Neural operators
[…]
@@ 3.1.1 Fourier neural operator
The FFT operation in Eq. (16) transforms the temporal sequence from the time domain to the frequency domain, where each frequency mode captures different temporal scales of the loading history. The learnable complex-valued filter $R_{\ell}$ acts as a frequency-selective operator, enabling the model to emphasize specific temporal patterns based on their relevance to the constitutive response. Importantly, the frequency-domain representation makes the learned operator independent of the temporal discretization used during training. Since the learned parameters index physical wavenumbers rather than discrete grid points, the same operator applies consistently across different sampling resolutions, enabling discretization-invariant predictions [43].
[…]
@@ 3.3 History-aware neural operator and its implementation
As shown in Eq. (20), HANO extends the standard neural operator framework by interleaving multiple standard Fourier layers with AEUF layers. Specifically, it leverages $\mathcal{L}^{\text{AEUF}}$ layers to further refine the latent representations produced by the preceding Fourier layers, through coupling the global spectral context with attention-guided, multiscale detail recovery, as discussed in Section 3.2. The embedded self-attention module dynamically re-weights the strain–stress history, allowing the model to selectively focus on the most relevant segments for next-step prediction. Within each $\mathcal{L}^{\text{AEUF}}$ block, a U-Net pathway recovers high-frequency and localized features that may otherwise be attenuated by the truncated Fourier basis, while a parallel Fourier branch retains coarse-scale global trends. This combination of U-Net, Fourier, and linear branches enables the AEUF layer to effectively fuse low- and high-frequency information in a single pass. By stacking multiple AEUF layers, HANO progressively enhances its capacity to model both long-range dependencies and localized transitions, leading to substantial improvements in path-dependent stress prediction, as demonstrated in Section 5.3. This HANO architecture serves as the default framework throughout this work and corresponds to the HANO3 variant discussed in Section 5.2.
@@ 3.3.1 Data structure of HANO for constitutive modeling
When the available stress-strain history is shorter than the required input length k (e.g., when loading begins from the virgin undeformed state), we employ zero-padding strategies to maintain consistent input dimensions. Specifically, missing historical data points are padded with zero stress-strain pairs, allowing the model to process loading histories of varying lengths while preserving the temporal structure of the input sequence.
@@
Remark 3.3 (Hidden variable independence): Different from many existing data-driven approaches that rely on reconstructing (physical or nominal) hidden variables encode loading history information [22, 46, 32], HANO directly leverages available strain–stress histories as input. This end-to-end formulation allows HANO to operate as a hidden-state-free model while capable of capturing the underlying dynamics of internal variable (ISV) evolution without explicitly modeling them. A key practical advantage of this autoregressive formulation is that it relies on observable quantities rather than latent, non-physical hidden states. If a material is under an unknown pre-stressed state, one can apply $k$ small controlled loading increments and record the corresponding stress–strain responses to initialize the model. This procedure can be carried out through controlled experiments on small specimens or, when the data-driven constitutive model is embedded within a finite element solver, through a few trial load steps at the integration point level. In this way, HANO achieves consistent and robust predictions from any arbitrary point along the loading path. This design naturally supports autoregressive forecasting, as further demonstrated in 4.2, where the model accurately predicts path-dependent material behavior even when initialized mid-trajectory.
@@ 3.3.2 Model prediction & loss function
where $n$ denotes the current step, and $k$ denotes the chosen window length. For multi‑step forecasting, we iterate (22) recursively: after predicting $\hat{\bm{\sigma}}_{n+1}$, we update the strain $\bm{\varepsilon}_{n+1}=\bm{\varepsilon}_{n}+\Delta\bm{\varepsilon}_{n+1}$, slide the window forward by discarding the oldest pair $(\bm{\varepsilon}_{n-k+1},\bm{\sigma}_{n-k+1})$ and appending the newly formed pair $(\bm{\varepsilon}_{n+1},\hat{\bm{\sigma}}_{n+1})$, and then supply the next increment $\Delta\bm{\varepsilon}_{n+2}$. In short, we set $n\rightarrow n+1$ in Eq. (22) in the next step. Repeating this procedure until the desired prediction length yields the complete stress trajectory. We note that, for forecasting in test cases, only the data within the initial window are required.
@@
Following standard machine learning practice [21], all strain and stress components are normalized to zero mean and unit variance using the mean and standard deviation computed from the training dataset. The same normalization parameters are consistently applied to both training and test sets, with model predictions denormalized before evaluation.
@@ 3.3.3 Training scheme
This training approach serves dual practical purposes. In early epochs, we keep inputs nearly clean, allowing the network to learn the fundamental stress-strain relationship without interference. As training progresses, we increase noise levels while simultaneously implementing scheduled sampling to phase out teacher forcing. By coupling scheduled sampling with noise injection, the training scheme gradually exposes the network to similar imperfect inputs it will create during inference. As teacher forcing is phased out and controlled noises are added, the model learns to attenuate its own prediction drift while remaining resilient to measurement noise. This joint strategy yields materially more reliable long-horizon stress forecasts, because the network is conditioned to operate under realistic inference conditions well. Representative training and validation loss curves demonstrating the effectiveness of this training strategy are provided in  E.
@@ 4.1 Data preparation & problem setting
The training and testing datasets used in this verification example are derived from the one-dimensional elastoplastic model with kinematic hardening, characterized by the Young’s modulus $E=200\,\text{GPa}$, the yield stress $\sigma_{y}=0.2\,\text{GPa}$, the hardening modulus $H=20\,\text{GPa}$. Details of the constitutive model and its implementation are provided in A. Starting from a stress‐free state (referred to as the undeformed configuration), we generate 1,000 distinct loading histories, each comprising two complete loading–unloading cycles and uniformly sampled with 100 increments (i.e., $T_{j}\equiv 100,\forall j$). The loading and unloading strain amplitudes $e_{\text{load}}$ and $e_{\text{unload}}$ are randomly sampled from the interval [0.008, 0.015][0.008,\,0.015] and [0.003, 0.007][0.003,\,0.007], respectively. Unless otherwise stated, all stress‐related quantities in this section are expressed in gigapascals (GPa).
@@
The effectiveness of the proposed HANO model was evaluated through a few benchmark tests against two variants of RNN-based models:
  • 1.
    RNN1: The original RNN model from Gorji et al. [22], designated as RNN1, which exclusively uses $k$ strain history data as input;
  • 2.
    RNN2: A modified RNN model, designated as RNN2, whose input layer is revised to receive the same concatenated strain–stress history window and forthcoming strain increment as the HANO model.
This systematic comparison enables us to isolate two key effects: (1) RNN1 vs RNN2 comparison isolates the impact of input representation and prediction strategy, where RNN2 uses identical input structure (strain-stress history + increment) as HANO and conducts prediction autoregressively, and (2) RNN2 vs HANO comparison isolates the architectural differences between finite-dimensional recurrent networks and operator learning frameworks under identical training protocols and input configurations.
All data-driven models (HANO and RNNs) are trained with the identical protocol described in Section 3.3.3. Table 1 summarizes the dataset configurations used for all numerical studies in this section based on the 1D elastoplastic material, while varying the training and testing configurations to evaluate different aspects of model performance.
Figure # Description Training Dataset Test Dataset
Figure 3 (Section 4.2.1) Evaluation of initial state sensitivity (comparing RNN1, RNN2, and HANO) 800 histories with kinematic hardening, 100 increments per cycle, zero-stress initialization (Default training dataset described in Section 4.1) Single case from Testset II with $\sim$45% of initial loading history removed, prediction from pre-stressed state ($\sigma\neq 0$, $\varepsilon^{p}\neq 0$)
Figure 4 (Section 4.3.2) Evaluation of resolution sensitivity in test dataset (comparing RNN2 and HANO) Same configuration as above Different loading paths from Testset I evaluated at multiple discretizations: 60, 70, 80, 90, 100, 120, 130, 140, 150 increments per cycle
Figure 5 (Section 4.3.2) Evaluation of discretization sensitivity: Prediction of stress–strain trajectories in representative test cases at different resolutions (comparing RNN2 and HANO) Same configuration as above Single representative test case evaluated at eight different resolutions (60-150 steps)
Figure 7 (Section 4.3.3) Evaluation of discretization invariance: HANO trained on variable-resolution data and tested across different resolutions Variable-resolution training: 1000 cases with randomly distributed resolutions (50-150 steps per cycle) Four test cases with unseen temporal resolutions: 57, 83, 112, and 135 increments per cycle
Figure 8 (Section 4.4) Evaluation of extrapolation capability beyond training scope Same as Figure 3 (Default training dataset described in Section 4.1) Extended 5-cycle loading sequences for extrapolation assessment
Figure 9 (Section 4.5) Evaluation of robustness to noisy input data Same configuration as above Gaussian noise added to the test inputs
Table 1: Summary of training and test dataset configurations used in each experiment presented in Section 4, based on the 1D elastoplastic material model introduced in Section 4.1.